OpenStudy (anonymous):

what is the lowest term of 24/30?

OpenStudy (shamil98):

24/30 they can both be divided by two. 12/15, this can be divided three 4/5, is your simplest form.

OpenStudy (shamil98):

you could've also just divided it by 6 (24/6) / (30/6) = 4/5

OpenStudy (the_fizicx99):

Both have a common multiple of 2 so 4/5 is my guess?

OpenStudy (anonymous):

like this...\[\frac{ 24 }{ 30 }=\frac{ 4\cdot \cancel{6} }{ 5\cdot \cancel{6} }=\frac{ 4 }{ 5 }\]